# Do Current Language Models Support Code Intelligence for R Programming Language?

### v2

@@ 4.3.2. Method Name Prediction
Table 7 represents the results for method name prediction in intra- and cross-project setting. The results highlight the consistent trend observed in prior studies involving intra and cross-project setting (Zhang et al., 2016b). Models generally perform better in intra-project settings compared to cross-project settings. This pattern is evident across all four models when applied to the R programming datasets.
@@ 4.4. RQ4: Code-PLM Performance When Separating R Style Datasets
In this section, we examine the performance of four baseline models fine-tuned on three distinct R programming datasets for two tasks of code summarization and name prediction, as shown in Figure 2. The goal is to evaluate how different styles of R and varying training setups impact the models’ ability to accurately summarize code and predict method names. We also examined the statistic significance of the difference between RCombine and two styles RTidy and RBase. Since the population size is not the same across RCombine, RTidy, and RBase we take random samples from RCombine to match the size of other two.
@@
  • For code summarization under monolingual training, the differences in performance are statistic significant for CodeBERT and CodeT5, the differences are only significant between RCombine and RBase for GraphCodeBERT and UniXCoder. When looking at name prediction under monolingual fine-tuning, the difference in performance is only statistic significant between RCombine and RBase for UniXCoder.
@@
  • For code summarization under multi-lingual fine-tuning, the differences in performance are statistic significant for all models except CodeT5, where the performance dropped 5% and 9.4% for RTidy and RBase, respectively. For name prediction under multi-lingual fine-tuning, similar trends were found, where the difference in performance is only statistic significant for one model, which is GraphCodeBERT.
@@
  • For code summarization under cross-project training, the difference in performance is statistic significant between RCombine and RTidy for CodeBERT, GraphCodeBERT and UniXCoder. The differences are more significant under intra-project training, where all of them are statistic significant.
@@
  • For method name prediction using cross-project training, the difference in performance is significant for CodeBERT and UniXCoder. It is only significant for GraphCodeBERT and CodeT5 between RCombine and RBase and RCombine and RTidy, respectively. Unlike code summarization, the differences are less significant under intra-project training, where all of them are not statistically significant different.
@@
In the code summarization task for the monolingual setting, the results indicate that models trained on the Base R dataset outperform those trained on the Tidy-verse and the combined datasets. The Base R models produced the highest scores for all baseline models except GraphCodeBERT, indicating that these models handle Base R syntax more effectively when summarizing code. The improvement in performance for Base R could be a benefit from a syntax structure that is simpler and closer to other programming languages. Tidy-verse-based models, while still performing relatively well, did not match the performance of RBase models, suggesting that the abstract and functional style of Tidy-verse poses additional challenges for the models during fine-tuning phase. The combined dataset, which merges both Base R and Tidy-verse, resulted in the lowest performance. This outcome implies that combining two distinct paradigms introduce complexities that interfere with the model’s ability to generate concise and accurate summaries. The performance shows that the two styles introduce conflicting patterns to the models and make the models struggle to reconcile.
@@
In contrast, in the name prediction task, the performance differences between the datasets are minimal. Whether the models were trained on Base R, Tidy-verse, or the combined dataset, the performance remained relatively similar. Previous work (Ahmed and Devanbu, 2022) shows that identifier plays the most important role in code tasks. This consistency indicates that method name prediction is less sensitive to the specific R style used, but more rely on the identifiers to capture the key information. Developers would use method names across Base R and Tidy-verse that share enough similarities to enable the models to generalize effectively across these paradigms. The combined dataset, although slightly weaker, did not exhibit the performance drop seen in code summarization. The consistency across different R programming styles reinforces the observation that name prediction is less dependent on variations in R programming styles, but mainly dependent on the identifiers.
@@
When shifting to the multilingual setup, the performance for both tasks showed a various degrees of decline. In the code summarization task, the same pattern observed in the monolingual setup persisted: models trained on Base R achieved the highest scores for most baseline models, while the combined dataset performed the worst. Contradicting previous research (Ahmed and Devanbu, 2022), the multilingual environment did not improve the performance of the models, it introduces additional complexity, as the models must handle syntax and structures from multiple programming languages that has different syntax and structure in many ways. This added burden likely impedes their ability to focus on the nuanced differences between Base R and Tidy-verse, making it more difficult to generate accurate code summaries, especially when trained on a mix of both paradigms.
@@ 4.5. RQ5: LLM Results
Table 8. Pass@k on HumanEval for selected LLMs.
@@ 4.6. Human evaluation
Table 9 shows the human evaluation scores received by each model for code summarization task. Each model is evaluated from three aspects: informativeness, relevance, and fluency.
@@
Table 10 shows the human evaluation scores for the baseline models. Semantic similarity measures how well the predicted names capture the meaning and purpose of the code elements they are naming. It is a crucial aspect of name prediction, as semantically appropriate names significantly enhance code readability and maintainability. CodeT5 stands out as the top performer with a score of $2.83$, which is notably higher than the ground truth score of $2.66$. This is an intriguing result, suggesting that CodeT5 is capable of generating names that human evaluators find even more semantically appropriate than the original names in some cases. This could indicate that CodeT5 has developed a strong understanding of code semantics and naming conventions, possibly even surpassing human performance in certain contexts. On the other hand, the low score received by the ground truth also indicates that some ground truth answer would make more sense when given the background information of the project.
GraphCodeBERT follows with a score of $2.66$, which interestingly matches the ground truth score exactly. This suggests that GraphCodeBERT’s name predictions are, on average, as semantically similar to the code’s purpose as the original names. CodeBERT comes in third with a score of $2.53$, which is relatively close to the ground truth. While not matching the performance of CodeT5 or GraphCodeBERT, it still demonstrates a decent ability to generate semantically appropriate names. UniXCoder shows the lowest performance among the models with a score of $2.27$, which is noticeably below the ground truth and other models. This suggests that UniXCoder may have more difficulty capturing the semantic nuances required for effective name prediction in this context. The relative performance of these models in name prediction aligns somewhat with their performance in code summarization (from the Table 10), with CodeT5 leading in both tasks. This consistency suggests that the underlying capabilities that make a model good at summarizing code also contribute to its ability to predict semantically appropriate names.
@@ 5. Discussions
In this section, we examine several factors influencing the observed performance in our study, particularly in the context of R programming. Key points include the choice of models employed, the impact of dataset, the models’ generalization capabilities, and the effectiveness of different training strategies.
@@ 5.1. R as a Low Resource Language
Figure 3. number of available projects on GitHub. Result is obtained by requesting GitHub through REST API with same query statements except programming language
From Figure 3, JavaScript, Python, and Java have approximately 10 times more GitHub repositories than R. Moreover, we can see that the available R projects is half of the size compare to Ruby, which is considered as low-resource language in the CodeSearchNet (Husain et al., 2019) dataset. The relatively low number of R packages on GitHub compared to other popular programming languages presents significant challenges for training effective language models and LLMs focused on R.
The smaller dataset also poses difficulties for transfer learning and fine-tuning approaches. We can see from Table 4 and Table 5, multi-lingual fine-tuning would not increase the performance even this approach has been showed to be effective in previous studies (Ahmed and Devanbu, 2022).
@@
CodeT5 Base: Converts a numeric matrix to a data.frame
@@
The dichotomy between Tidy-verse and Base R in the R programming presents an interesting challenge for language models and LLMs trained on R code. This division could significantly impact model performance and introduces unique considerations for models working with R. Tidy-verse, a collection of popular R packages designed for data science, introduces a distinct syntax and philosophy that differs from Base R. This split in coding styles and approaches within the R community can affect language models in several ways. Models trained on a dataset that includes code from both R styles, may struggle to maintain consistency in their outputs, potentially mixing idioms. This is more challenging specifically interfere code generation by mixing both paradigms. This could lead to confusion for users expecting pure Tidy-verse or Base R solutions. For example, the summaries generated by CodeLlama2 for both Tidy and Base shows a good understanding of the code’s functionality but the Tidy version shows less information about the function compared to the Base version.
@@
Listings 3 and  4 are the Base and Tidy version of the same function. The function processes a numeric matrix by first validating the input to ensure it meets the minimum criteria. It transposes the matrix, converts it into a data frame, and renames the columns systematically as $y_{1},y_{2},\dots,y_{n}$. Optionally, it filters duplicated rows based on user specification, ensures that the data is structured and ready for further analysis or modeling. The Base R implementation, as shown in the Listing 3 example, follows a more traditional, imperative style, relying on Base R functions like as.data.frame(), t(), and filterDuplicated(). In contrast, the Tidy-verse version, listing 4, embraces a more expressive, declarative approach, utilizing as_tibble(), rename_with(), and distinct() from the Tidy-verse packages.
@@
We provided a list of summarizations from each model in Table 11, where they are asked to summarize a Base and Tidy version of the same function shown in Listings 3 and  4. The pre-trained language models are trained for code summarization and the LLMs are doing this task under a zero-shot setup. The performance of the analyzed models reflects their varying levels of exposure to and familiarity with these R programming idioms. CodeBERT ,GraphCodeBERT and CodeT5, demonstrate a stronger grasp of the underlying concepts when compared to UniXcoder, as evidenced by their precise summaries for both the Base R and Tidy-verse versions. However, the summaries provided by those models are very short in general, missing key information about the function’s underlying task. For example, only CodeT5 in its summary for the Base version, motioned that the matrix need to be numeric. Moreover, the word “rename” is not accurate about the function’s functionality. In contrast, when we look at the summaries provided by the LLMs, one of the main difference is that LLM’s summaries are longer, and contain more information. For instance, CodeLlama provided information about the input and output of the function. Moreover, StarCoder provides summaries about what happened during the execution. On the other hand, StarCoder did not mention that the matrix is transposed into a dataframe when summarizing Tidy version, and missed the numeric part for Base version.
@@
Furthermore, all models’ ability to capture project-level context and constraints, such as domain-specific requirements, implicit dependencies, and established coding conventions, appears to be an area for potential improvement. All of the models missed the part about checkmate::assertMatrix, where it is a function from a larger context. The summaries generated by the models, while accurate in conveying the general purpose of the functions, lack the nuanced understanding of how these code snippets would integrate within the larger context of a real-world R project. This is an existing challenge for other programming languages, as shown in previous studies in the existance of unfamiliar libraries or private repositories (Zhou et al., 2022; Jimenez et al., 2023). Previous work (Yun et al., 2024) also shows that including context information would increase the performance for code summarization task.
@@
Moreover, from the examples provided in listing 3 and listing 4, we can see that the prevalence of Tidy-verse in modern R programming might skew model performance. If the training data heavily favors Tidy-verse code, the model may underperform when dealing with Base R tasks, and vice versa. This imbalance could result in biased code suggestions or completions, potentially steering users towards one paradigm over the other unintentionally, especially when Tidy-verse is heavily used in some domain while barely used in others. The model’s ability to understand and generate idiomatic R code could vary significantly depending on whether it is working with Tidy-verse or Base R constructs.

### v3

@@ 4.3.2. Method Name Prediction
Table 7 represents the results for method name prediction in intra- and cross-project settings. The results highlight the consistent trend observed in prior studies involving intra and cross-project settings (Zhang et al., 2016b). Models generally perform better in intra-project settings compared to cross-project settings. This pattern is evident across all four models when applied to the R programming datasets.
@@ 4.4. RQ4: Code-PLM Performance When Separating R Style Datasets
In this section, we examine the performance of four baseline models fine-tuned on three distinct R programming datasets for two tasks of code summarization and name prediction, as shown in Figure 2. The goal is to evaluate how different styles of R and varying training setups impact the models’ ability to accurately summarize code and predict method names. We also examined the statistic significance of the difference between RCombine and the two styles, RTidy and RBase. Since the population size is not the same across RCombine, RTidy, and RBase, we take random samples from RCombine to match the size of the other two.
@@
  • For code summarization under monolingual training, the differences in performance are statistic significant for CodeBERT and CodeT5, the differences are only significant between RCombine and RBase for GraphCodeBERT and UniXCoder. When looking at name prediction under monolingual fine-tuning, the difference in performance is only statistically significant between RCombine and RBase for UniXCoder.
@@
  • For code summarization under multi-lingual fine-tuning, the differences in performance are statistically significant for all models except CodeT5, where the performance dropped 5% and 9.4% for RTidy and RBase, respectively. For name prediction under multi-lingual fine-tuning, similar trends were found, where the difference in performance is only statistically significant for one model, which is GraphCodeBERT.
@@
  • For code summarization under cross-project training, the difference in performance is statistically significant between RCombine and RTidy for CodeBERT, GraphCodeBERT, and UniXCoder. The differences are more significant under intra-project training, where all of them are statistically significant.
@@
  • For method name prediction using cross-project training, the difference in performance is significant for CodeBERT and UniXCoder. It is only significant for GraphCodeBERT and CodeT5 between RCombine and RBase and RCombine and RTidy, respectively. Unlike code summarization, the differences are less significant under intra-project training, where all of them are not statistically significantly different.
@@
In the code summarization task for the monolingual setting, the results indicate that models trained on the Base R dataset outperform those trained on the Tidy-verse and the combined datasets. The Base R models produced the highest scores for all baseline models except GraphCodeBERT, indicating that these models handle Base R syntax more effectively when summarizing code. The improvement in performance for Base R could be a benefit from a syntax structure that is simpler and closer to other programming languages. Tidy-verse-based models, while still performing relatively well, did not match the performance of RBase models, suggesting that the abstract and functional style of Tidy-verse poses additional challenges for the models during the fine-tuning phase. The combined dataset, which merges both Base R and Tidy-verse, resulted in the lowest performance. This outcome implies that combining two distinct paradigms introduce complexities that interfere with the model’s ability to generate concise and accurate summaries. The performance shows that the two styles introduce conflicting patterns to the models and make the models struggle to reconcile.
@@
In contrast, in the name prediction task, the performance differences between the datasets are minimal. Whether the models were trained on Base R, Tidy-verse, or the combined dataset, the performance remained relatively similar. Previous work (Ahmed and Devanbu, 2022) shows that the identifier plays the most important role in code tasks. This consistency indicates that method name prediction is less sensitive to the specific R style used, but relies more on the identifiers to capture the key information. Developers would use method names across Base R and Tidy-verse that share enough similarities to enable the models to generalize effectively across these paradigms. The combined dataset, although slightly weaker, did not exhibit the performance drop seen in code summarization. The consistency across different R programming styles reinforces the observation that name prediction is less dependent on variations in R programming styles, but mainly dependent on the identifiers.
@@
When shifting to the multilingual setup, the performance for both tasks showed various degrees of decline. In the code summarization task, the same pattern observed in the monolingual setup persisted: models trained on Base R achieved the highest scores for most baseline models, while the combined dataset performed the worst. Contradicting previous research (Ahmed and Devanbu, 2022), the multilingual environment did not improve the performance of the models; it introduces additional complexity, as the models must handle syntax and structures from multiple programming languages that has different syntax and structure in many ways. This added burden likely impedes their ability to focus on the nuanced differences between Base R and Tidy-verse, making it more difficult to generate accurate code summaries, especially when trained on a mix of both paradigms.
@@
4.4.1. Model Performance by Code Length
Figure 3. Performance of baseline models on Tidy and Base across different code lengths in R under monolingual setups.
Figure 3 compares the performance of four baseline models—CodeBERT, GraphCodeBERT, UnixCoder, and CodeT5—across code segments of varying lengths in R, while contrasting the Tidy-verse and Base versions of the dataset for code summarization. A clear trend emerges in which shorter code snippets (for instance, those under 50 tokens) achieve higher BLEU scores. This outcome is not surprising, as shorter snippets frequently lack the level of complexity found in longer blocks of code, making it easier for the models to learn mappings between source code and summary. Fewer tokens also reduce the chance of encountering convoluted control flow or deeply nested function calls, leading to more straightforward patterns that the models can accurately capture.
As the code length increases beyond around 150 or 200 tokens, performance begins to decrease, reflecting the additional challenges posed by complex logic, multiple layers of function calls, and generally higher token diversity. When code is longer, the model must handle a broader context and maintain a more extensive memory of previous tokens to accurately capture meaning. The chart further shows that, within the same length range, the Tidy-verse data often confers a noticeable advantage over Base data (except for CodeBERT). Tidy-verse data removes inconsistencies such as irregular spacing or extraneous comments, thereby reducing noise and helping the model focus on core functional elements of the code. This benefit becomes even more significant for longer snippets, where seemingly minor inconsistencies can accumulate and hinder the model’s ability to process and summarize effectively. Additionally, Tidy-verse is focused on data science, providing a consistent and intuitive syntax (such as the pipe operator and a series of well-defined functions) that makes data cleaning, transformation, and visualization more efficient and readable [^12] https://tidyverse.tidyverse.org/articles/paper.html. This design approach is similar to the usage of data science libraries like pandas in Python, as both emphasize code readability and coherent operations. Therefore, the syntax and usage could be more specific compared to Base R, and this difference might affect the results.
@@ 4.5. RQ5: LLM Results
Table 8. Pass@k on HumanEval for selected LLMs. The temperature is set to 0.8.
@@
few-shot-example Model Zero Shot 1-Shot 3-Shot 5-Shot 10-Shot
Simple R function codegen-2B-multi 4.3 4.3 5.6 5.6 0.6
CodeLlama-7B 21.1 17.4 18.6 11.8 0.0
StarCoder2-7B 20.5 19.3 22.4 22.4 9.9
BM25 codegen-2B-multi 4.3 1.24 0.0 0.0 0.0
CodeLlama-7B 21.1 7.45 0.0 0.0 0.0
StarCoder2-7B 20.5 7.45 0.0 0.0 0.0
Embedding codegen-2B-multi 4.3 1.86 1.86 0.62 0.0
CodeLlama-7B 21.1 10.56 12.42 3.73 0.0
StarCoder2-7B 20.5 15.53 11.80 8.70 0.0
Table 9. Model Performance at various shot counts applied to R. Results represent pass@1. The temperature is set to 0, and we follow a greedy approach.
Previous research shows that adding examples as context to the models improves their performance for code generation (Nashid et al., 2023). Therefore, we conducted additional experiments, in which we provided the models with additional context, being 1, 3, 5, and 10 samples. We applied three settings to add additional context. In the first setting, the examples are chosen from R tutorials and/or written by an R expert. These are simple examples showing, for example, how to import a csv file or read a column from a tabular dataset. Then these examples are used as the additional context to the model. In the second setting, we used the BM25 algorithm to choose the top $K$ similar samples from our collected dataset, where $k\in\{1,3,5,10\}$ and chose them as the context in the few-shot setting. In the third setting, we calculated the cosine similarity between the embeddings of our collected dataset and the question prompt, and chose the top $K$ similar samples from the dataset as the context in the few-shot setting.
Table 9 presents a performance comparison between three few-shot strategies—reusing the same, simple R function as the examples for all queries, BM25, and by cosine similarity between the embeddings. In the first setting, each of the three models produces higher scores compared to using BM25, reaching as high as $22.4$ (3-shot using StarCoder2-7B). This indicates that even a single, generic example can still provide enough information for the models to solve the tasks in the HumanEval setting. By contrast, the BM25-based approach shows consistently low performance, with most scores dropping to zero for all but the zero-shot column. Two main reasons explain this performance gap. First, the code snippets retrieved via BM25, though superficially related to the query, often do not match the HumanEval tasks in either style or goal, meaning the models do not benefit much from these examples. Second, the retrieved examples can be lengthy, so the models fail to utilize the content effectively. We see the same trend for the embedding-based approach, where the model shows consistently decreased performance. However, the embedding-based approach outperformed the BM25-based approach, where CodeLlama-7B and StarCoder2-7B achieved measurable improvements at different shot settings, for example, achieving $11.80$ for CodeLlama-7B and $12.42$ for StarCoder2-7B in 3-shot compared to 00 of BM25. Embedding methods appear more adept at selecting relevant context than keyword-based BM25, showing that vector similarities capture semantic relationships beyond surface-level token matches. Compared to the zero-shot approach, providing 3 examples in the prompt helps improve the scores slightly. However, the Simple R function has similar scores with the embedding setting. Simple R function provides a simpler and more focused demonstration within the context window of the model, which helps generate the code for the test dataset. Another interesting observation is that adding more examples does not improve the performance, as shown by the performance drop for 5-shot for CodeLlama-7B and all models in the 10-shot setting. We attribute this behavior to the lengthy input provided to the model.
@@ 4.6. Human evaluation
Table 10 shows the human evaluation scores received by each model for the code summarization task. Each model is evaluated from three aspects: informativeness, relevance, and fluency.
@@
Table 11 shows the human evaluation scores for the baseline models. Semantic similarity measures how well the predicted names capture the meaning and purpose of the code elements they are naming. It is a crucial aspect of name prediction, as semantically appropriate names significantly enhance code readability and maintainability. CodeT5 stands out as the top performer with a score of $2.83$, which is notably higher than the ground truth score of $2.66$. This is an intriguing result, suggesting that CodeT5 is capable of generating names that human evaluators find even more semantically appropriate than the original names in some cases. This could indicate that CodeT5 has developed a strong understanding of code semantics and naming conventions, possibly even surpassing human performance in certain contexts. On the other hand, the low score received by the ground truth also indicates that some ground truth answers would make more sense when given the background information of the project.
GraphCodeBERT follows with a score of $2.66$, which interestingly matches the ground truth score exactly. This suggests that GraphCodeBERT’s name predictions are, on average, as semantically similar to the code’s purpose as the original names. CodeBERT comes in third with a score of $2.53$, which is relatively close to the ground truth. While not matching the performance of CodeT5 or GraphCodeBERT, it still demonstrates a decent ability to generate semantically appropriate names. UniXCoder shows the lowest performance among the models with a score of $2.27$, which is noticeably below the ground truth and other models. This suggests that UniXCoder may have more difficulty capturing the semantic nuances required for effective name prediction in this context. The relative performance of these models in name prediction aligns somewhat with their performance in code summarization (from Table 11), with CodeT5 leading in both tasks. This consistency suggests that the underlying capabilities that make a model good at summarizing code also contribute to its ability to predict semantically appropriate names.
@@ 5. Discussions
In this section, we examine several factors influencing the observed performance in our study, particularly in the context of R programming. Key points include the choice of models employed, the impact of the dataset, the models’ generalization capabilities, and the effectiveness of different training strategies.
@@ 5.1. R as a Low Resource Language
Figure 4. number of available projects on GitHub. The result is obtained by requesting GitHub through the REST API with the same query statements, except programming language.
From Figure 4, JavaScript, Python, and Java have approximately 10 times more GitHub repositories than R. Moreover, we can see that the available R projects are half of the size compare to Ruby, which is considered a low-resource language in the CodeSearchNet (Husain et al., 2019) dataset. The relatively low number of R packages on GitHub compared to other popular programming languages presents significant challenges for training effective language models and LLMs focused on R.
The smaller dataset also poses difficulties for transfer learning and fine-tuning approaches. We can see from Table 4 and Table 5, multi-lingual fine-tuning would not increase the performance, even though this approach has been showed to be effective in previous studies (Ahmed and Devanbu, 2022).
@@
CodeT5 Base: Converts a numeric matrix to a dataframe
@@
The dichotomy between Tidy-verse and Base R in the R programming presents an interesting challenge for language models and LLMs trained on R code. This division could significantly impact model performance and introduce unique considerations for models working with R. Tidy-verse, a collection of popular R packages designed for data science, introduces a distinct syntax and philosophy that differs from Base R. This split in coding styles and approaches within the R community can affect language models in several ways. Models trained on a dataset that includes code from both R styles, may struggle to maintain consistency in their outputs, potentially mixing idioms. This is more challenging, specifically interfering with code generation by mixing both paradigms. This could lead to confusion for users expecting pure Tidy-verse or Base R solutions. For example, the summaries generated by CodeLlama2 for both Tidy and Base show a good understanding of the code’s functionality, but the Tidy version shows less information about the function compared to the Base version.
@@
Listings 3 and  4 are the Base and Tidy versions of the same function. The function processes a numeric matrix by first validating the input to ensure it meets the minimum criteria. It transposes the matrix, converts it into a data frame, and renames the columns systematically as $y_{1},y_{2},\dots,y_{n}$. Optionally, it filters duplicated rows based on user specification, ensuring that the data is structured and ready for further analysis or modeling. The Base R implementation, as shown in the Listing 3 example, follows a more traditional, imperative style, relying on Base R functions like as.data.frame(), t(), and filterDuplicated(). In contrast, the Tidy-verse version, listing 4, embraces a more expressive, declarative approach, utilizing as_tibble(), rename_with(), and distinct() from the Tidy-verse packages.
@@
We provided a list of summarizations from each model in Table 12, where they are asked to summarize a Base and Tidy version of the same function shown in Listings 3 and  4. The pre-trained language models are trained for code summarization, and the LLMs are doing this task under a zero-shot setup. The performance of the analyzed models reflects their varying levels of exposure to and familiarity with these R programming idioms. CodeBERT, GraphCodeBERT and CodeT5 demonstrate a stronger grasp of the underlying concepts when compared to UniXCoder, as evidenced by their precise summaries for both the Base R and Tidy-verse versions. However, the summaries provided by those models are very short in general, missing key information about the function’s underlying task. For example, only CodeT5 in its summary for the Base version mentioned that the matrix need to be numeric. Moreover, the word “rename” is not accurate about the function’s functionality. In contrast, when we look at the summaries provided by the LLMs, one of the main difference is that LLM’s summaries are longer, and contain more information. For instance, CodeLlama provided information about the input and output of the function. Moreover, StarCoder provides summaries about what happened during the execution. On the other hand, StarCoder did not mention that the matrix is transposed into a dataframe when summarizing Tidy version, and missed the numeric part for Base version.
@@
Furthermore, all models’ ability to capture project-level context and constraints, such as domain-specific requirements, implicit dependencies, and established coding conventions, appears to be an area for potential improvement. All of the models missed the part about checkmate::assertMatrix, where it is a function from a larger context. The summaries generated by the models, while accurate in conveying the general purpose of the functions, lack the nuanced understanding of how these code snippets would integrate within the larger context of a real-world R project. This is an existing challenge for other programming languages, as shown in previous studies in the existance of unfamiliar libraries or private repositories (Zhou et al., 2022; Jimenez et al., 2023). Previous work (Yun et al., 2024) also shows that including context information would increase the performance for the code summarization task.
@@
Moreover, from the examples provided in listing 3 and listing 4, we can see that the prevalence of Tidy-verse in modern R programming might skew model performance. If the training data heavily favors Tidy-verse code, the model may underperform when dealing with Base R tasks, and vice versa. This imbalance could result in biased code suggestions or completions, potentially steering users towards one paradigm over the other unintentionally, especially when Tidy-verse is heavily used in some domains while barely used in others. The model’s ability to understand and generate idiomatic R code could vary significantly depending on whether it is working with Tidy-verse or Base R constructs.